# Byam: Fixing Breaking Dependency Updates with Large Language Models

### v1

@@ Abstract
Application Programming Interfaces (APIs) facilitate the integration of third-party dependencies within the code of client applications. However, changes to an API, such as deprecation, modification of parameter names or types, or complete replacement with a new API, can break existing client code. These changes are called breaking dependency updates; It is often tedious for API users to identify the cause of these breaks and update their code accordingly. In this paper, we explore the use of Large Language Models (LLMs) to automate client code updates in response to breaking dependency updates. We evaluate our approach on the BUMP dataset, a benchmark for breaking dependency updates in Java projects. Our approach leverages LLMs with advanced prompts, including information from the build process and from the breaking dependency analysis. We assess effectiveness at three granularity levels: at the build level, the file level, and the individual compilation error level. We experiment with five LLMs: Google Gemini-2.0 Flash, OpenAI GPT4o-mini, OpenAI o3-mini, Alibaba Qwen2.5-32b-instruct, and DeepSeek V3. Our results show that LLMs can automatically repair breaking updates. Among the considered models, OpenAI’s o3-mini is the best, able to completely fix 27% of the builds when using prompts that include contextual information such as the buggy line, API differences, error messages, and step-by-step reasoning instructions. Also, it fixes 78% of the individual compilation errors. Overall, our findings demonstrate the potential for LLMs to fix compilation errors due to breaking dependency updates, supporting developers in their efforts to stay up-to-date with changes in their dependencies.
Keywords:
Breaking Dependency Update Breaking Changes API evolution LLMs
@@ 1 Introduction
Modern software development heavily relies on third-party software libraries to improve developers’ productivity and time-to-market. The set of libraries used by a client project are referred to as the dependencies of that project. Developers access a library’s functionality through its provided Application Programming Interfaces (APIs). These APIs evolve over time to introduce new functionality, fix bugs, or address security vulnerabilities (Lamothe et al., 2021).
The evolution of library APIs can cause client code to fail. This is due to modifications such as changes in a parameter name or type, method deprecation, or even complete removal of classes and modules  (Lamothe et al., 2021; Reyes et al., 2024b ; Reyes et al., 2024a ). Failures that are due to API updates are referred to as breaking dependency updates. When a new version of a dependency results in a breaking dependency update, the developers using the library (i.e., the client developers) have to figure out how to update their code to fix the errors they face. This is known to be a tedious and time consuming maintenance task (Zhang et al., 2022; Ochoa et al., 2022).
@@
Breaking dependency updates are common in software projects (Brito et al., 2018b ), and often discourage developers from upgrading their dependencies (Dietrich et al., 2019; Ochoa et al., 2022; Venturini et al., 2023). However, keeping dependencies up to date is crucial for security, as outdated libraries may contain vulnerabilities  (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020; Larios Vargas et al., 2020; Salza et al., 2018).
Most prior work on breaking dependency updates focused on analyzing their frequency, causes, and identifying breaking changes in library code (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020; Brito et al., 2018a ; Mujahid et al., 2020). There is scarce work on automating client code fixes. Xing et al. (Xing and Stroulia, 2007) proposed rule-based program transformation techniques, limited to predefined patterns and small adaptation examples. Other approaches improved rule synthesis by incorporating additional library usage examples (Brito et al., 2018a ) or recommendations for API method replacements based on source code analysis (Dagenais and Robillard, 2009). These approaches struggle to handle the widely diverse causes of breaking updates, such as method signature alterations, and removed classes. This is the problem we address in this paper.
In this work, we explore the use of Large Language Models (LLMs) to automate the generation of a patch to fix a breaking dependency update. LLMs are neural networks trained on large datasets of text and code (Vaswani, 2017; Chang et al., 2024). They have shown great effectiveness in software engineering tasks like code generation and bug fixing (Nam et al., 2024; Yin et al., 2024; Hidvégi et al., 2024).
We design and evaluate Byam, a pipeline that leverages LLMs to update client code in response to breaking dependency updates. We focus on breaking dependency updates that lead to compilation errors, which are the most common types of causes for breakage (Reyes et al., 2024b ; Reyes et al., 2024a ). Byam analyzes compilation errors caused by a dependency update, formulates prompts that incorporate relevant build and dependency information, and queries an LLM to generate updated code. We assess the effectiveness of Byam using the BUMP dataset  (Reyes et al., 2024b ), which provides reproducible cases of breaking dependency updates in Java projects. Our experiments systematically evaluate different LLMs, prompt design strategies, and information granularity levels.
We evaluate Byam with five different LLMs: Google Gemini-2.0 Flash, OpenAI GPT4o-mini, OpenAI o3-mini, Alibaba Qwen2.5-32b-instruct, and DeepSeek V3. We assess its effectiveness at three levels of granularity: build, file, and individual error. At the build level, we measure the number of fully fixed builds, those still failing due to compilation errors, and cases where compilation was resolved but test failures emerged. At the file level, we track the number of files with compilation errors before and after the fixes, identifying which files were successfully fixed and which ones still include compilation errors. Finally, at the individual error level, we analyze the number of original compilation errors per file, comparing them before and after the LLM-generated fixes to determine fixed, unresolved, and newly introduced errors. To quantify the impact of the corrections generated by LLMs, we define a set of metrics to evaluate the performance of LLMs. We analyze the capability of the models to fully resolve the build process, file-level success rate, and performance in resolving individual errors. Since LLMs can also introduce new errors in the process, we measure the relative error fixed as the proportion of previously existing errors that were successfully resolved, while accounting for any new errors introduced after the code update.
Our results show that LLM can automatically repair breaking updates. OpenAI’s o3-mini model achieves the best performance with a build success rate of 27% when using the best prompt that integrates buggy line, information regarding the differences between the dependency versions (APIDiff), and structured reasoning techniques (CoT). At the error level, o3-mini fixes 79% of compilation failures. Overall, our experimental results demonstrate the potential for LLMs to reduce developer effort and time in managing breaking dependency updates, helping developers more effectively stay up-to-date with changes in their dependencies.
@@
    A new approach, called Byam, for fixing compilation errors from breaking dependency updates with LLMs. Its advanced prompting strategies relies on a comprehensive analysis of the breaking update problem space. In particular, we use API differencing of the breaking library and a Chains-of-thought reasoning command dedicated to breaking updates.
@@
    A large scale experiment with five different LLMs and eight prompts, over 103 breaking dependency updates. Our results show that o3-mini achieves the best outcomes when using a prompt that highlights the buggy line causing the compilation error(s), that provides information about the API differences between the old and updated dependency versions, and that includes Chains-of-thought instructions .
@@
2 Preliminaries
In this section, we define the key terminology used throughout this work.
@@ 2 Preliminaries
A breaking dependency update is a pair of commits for a project, consisting of a pre-breaking commit with a passing build and a breaking commit that updates the version of a single dependency, causing the build to fail.
@@
A breaking commit is the project commit where the only change is an update to a dependency version that leads to the project build failing.
Breaking dependency updates trigger different types of failures, such as compilation errors, test failures, dependency resolution failures, or lock failures (Reyes et al., 2024b ). In this work, we specifically focus on compilation failures, which are the most prevalent issue triggered by breaking dependency updates (Reyes et al., 2024a ).
3 Approach to Repairing Breaking Updates
@@ 3.2 Step 1: Extract Failed Build Information
This step isolates the parts of the code that are causing a compilation failure after the dependency update and identifies the API differences involved in the failure. We use this information to provide the LLM with various contextual data. In order to collect this information, we analyze the build output log after a breaking update. Byam extracts a list of files causing the compilation error(s) and the list of error messages for each file. Additionally, Byam analyzes the changes between both versions of the dependency.
Figure 1(c) shows an excerpt of a build output. Analyzing this output, we determine that the file /billy/billy-gin/src/main/java/com/premiumminds/billy/gin/services/impl/pdf/FO%PPDFTransformer.java is causing a compilation error. We also extract the error message, “no suitable method found for newInstance(no arguments), as well as the error location, at line 115, column 43. From the error location, we identify the specific constructs responsible for the compilation failure. In the previous example, the error is caused by a method newInstanc(). Next, Byam extracts the API differences (APIDiff) between the dependency versions to determine whether any changes in the new version are related to the construct that triggers the error. For this example, APIDiff reports that the signature of the method newInstance()has been changed in the new version 2.0 of the dependency. Figure 3 shows an excerpt of the APIDiff corresponding to the example in Figure 1, highlighting the change associated with the construct identified as the root cause of the breaking dependency update. This APIDiff information is later incorporated into the LLM prompts, as discussed in Subsection 3.3.
@@
$+$ NEW METHOD: PUBLIC(+) STATIC(+) org.apache.fop.apps.FopFactory newInstance(org.apache.fop.apps.FopFactoryConfig)
$+$ NEW METHOD: PUBLIC(+) STATIC(+) org.apache.fop.apps.FopFactory newInstance(java.io.File)
$+$ NEW EXCEPTION: org.xml.sax.SAXException
$+$ NEW EXCEPTION: java.io.IOException
$+$ NEW METHOD: PUBLIC(+) STATIC(+) org.apache.fop.apps.FopFactory newInstance(java.net.URI)
$+$ NEW METHOD: PUBLIC(+) STATIC(+) org.apache.fop.apps.FopFactory newInstance(java.net.URI, java.io.InputStream)
$+$ NEW EXCEPTION: org.xml.sax.SAXException
$+$ NEW EXCEPTION: java.io.IOException
@@ 3.3 Step 2: Use LLM to Generate Repairs
Buggy Line Inclusion Adds the specific line of code causing the compilation error.
@@
Based on the baseline, we evaluate specific combinations of prompt design choices rather than generating all possible variations, focusing on buggy lines, API version differences (APIDiff), and Chain of Thought (CoT) reasoning. Instead of an exhaustive exploration, we systematically test selected configurations that provide meaningful information on their impact on LLM performance. Our approach considers the incremental addition of each factor to a baseline prompt, ensuring a balanced assessment of each factor’s contributions. This approach examines two key dimensions: the prompting strategy (zero-shot vs. CoT) and the type of contextual information included in the prompt (error information, buggy line, APIDiff). This structured evaluation allows us to isolate the effects of each component while maintaining a manageable number of experimental setups.
Providing Buggy Line Information. Developers, when debugging, typically start by identifying the exact line of code causing the issue, as this helps narrow down the scope of the problem and focus efforts on the relevant section of the code. Including the buggy line in the prompt replicates this natural debugging process, potentially improving code understanding for the LLM (the compiler error message does not include the actual line by default). We include this information by extending the base prompt with the actual line of the client code responsible for the compilation error, as shown in Figure 5, after the error message.
@@
Figure 5: Adding Buggy Line to the Prompt. This helps the LLM to identify and focus on the problem to be fixed.
@@
To extract these differences, we rely on japicmp (Maven, 2022), a widely used tool to detect API changes in Java libraries. This tool identifies modifications such as changes in method signatures, parameter types, deprecated methods, or new functionality. For example, when we use japicmp to analyze the changes between versions 6.18.1 and 6.19.1 of the jasperreports dependency, it detects that the setLinewith method was removed from the new version of the dependency. We augment the prompt with that information, as in the example shown in Figure 6. Figure 7 shows the concrete example of APIDiff for jasperreports.
@@
$P_{4}$ Buggy Line + APIDiff
@@
$P_{8}$ CoT + Buggy Line + APIDiff

### v2

@@ Abstract
Application Programming Interfaces (APIs) facilitate the integration of third-party dependencies within the code of client applications. However, changes to an API, such as deprecation, modification of parameter names or types, or complete replacement with a new API, can break existing client code. These changes are called breaking dependency updates; It is often tedious for API users to identify the cause of these breaks and update their code accordingly. In this paper, we explore the use of Large Language Models (LLMs) to automate client code updates in response to breaking dependency updates. We evaluate our approach on the BUMP dataset, a benchmark for breaking dependency updates in Java projects. Our approach leverages LLMs with advanced prompts, including information from the build process and from the breaking dependency analysis. We assess effectiveness at three granularity levels: at the build level, the file level, and the individual compilation error level. We experiment with five LLMs: Google Gemini-2.0 Flash, OpenAI GPT4o-mini, OpenAI o3-mini, Alibaba Qwen2.5-32b-instruct, and DeepSeek V3. Our results show that LLMs can automatically repair breaking updates. Among the considered models, OpenAI’s o3-mini is the best, able to completely fix 27% of the builds when using prompts that include contextual information such as the erroneous line, API differences, error messages, and step-by-step reasoning instructions. Also, it fixes 78% of the individual compilation errors. Overall, our findings demonstrate the potential for LLMs to fix compilation errors due to breaking dependency updates, supporting developers in their efforts to stay up-to-date with changes in their dependencies.
@@ 1 Introduction
Modern software development heavily relies on third-party software libraries to improve developers’ productivity and time-to-market. The set of libraries used by a client project is referred to as the dependencies of that project. Developers access a library’s functionality through its provided Application Programming Interfaces (APIs). These APIs evolve over time to introduce new functionality, fix bugs, or address security vulnerabilities (Lamothe et al., 2021). While such evolution is necessary, it often comes at a cost: updates may introduce breaking changes that prevent client projects from compiling or running  (Lamothe et al., 2021; Reyes et al., 2024b ; Reyes et al., 2024a ). Failures resulting from API updates are referred to as breaking dependency updates. When a new version of a dependency results in a breaking dependency update, the developers using the library (i.e., the client developers) have to figure out how to update their code to fix the errors they face. This is known to be a tedious and time consuming maintenance task (Zhang et al., 2022; Ochoa et al., 2022).
To address this problem, we propose Byam, a pipeline that leverages Large Language Models (LLMs) to update client code in response to breaking dependency updates (Vaswani, 2017; Chang et al., 2024). We focus on breaking dependency updates that lead to compilation errors, which are the most common types of causes for breakage (Reyes et al., 2024b ; Reyes et al., 2024a ). Byam analyzes compilation errors caused by a dependency update, formulates prompts that incorporate relevant build and dependency information, and queries an LLM to generate updated code. We assess the effectiveness of Byam using the BUMP dataset  (Reyes et al., 2024b ), which provides reproducible cases of breaking dependency updates in Java projects. Our experiments systematically evaluate different LLMs, prompt design strategies, and information granularity levels.
We evaluate Byam with five different LLMs: Google Gemini-2.0 Flash, OpenAI GPT4o-mini, OpenAI o3-mini, Alibaba Qwen2.5-32b-instruct, and DeepSeek V3. We assess Byam’s effectiveness at three levels of granularity: build, file, and individual error. At the build level, we measure the number of fully fixed builds, those still failing due to compilation errors, and cases where compilation was resolved but test failures emerged. At the file level, we track the number of files with compilation errors before and after the fixes, identifying which files were successfully fixed and which ones still include compilation errors. Finally, at the individual error level, we analyze the number of original compilation errors per file, comparing them before and after the LLM-generated fixes to determine fixed, unresolved, and newly introduced errors. To quantify the impact of the corrections generated by LLMs, we define a set of metrics to evaluate the performance of LLMs. We analyze the capability of the models to fully resolve the build process, file-level success rate, and performance in resolving individual errors. Since LLMs can also introduce new errors in the process, we measure the relative error fixed as the proportion of previously existing errors that were successfully resolved, while accounting for any new errors introduced after the code update.
@@
Our results show that LLM can automatically repair breaking updates. OpenAI’s o3-mini model achieves the best performance with a build success rate of 27% when using the best prompt that integrates erroneous line, information regarding the differences between the dependency versions (APIDiff), and structured reasoning techniques (CoT). At the file level o3-mini fixes 41% of the originally erroneous files in unsuccessful builds and at the error level, o3-mini fixes 79% of compilation failures. Overall, our experimental results demonstrate the potential for LLMs to reduce developer effort and time in managing breaking dependency updates, helping developers more effectively stay up-to-date with changes in their dependencies.
@@
    An approach, called Byam, for fixing compilation errors from breaking dependency updates with LLMs. Its advanced prompting strategies relies on a comprehensive analysis of the breaking update problem space. In particular, we use API differencing of the breaking library and a Chains-of-thought reasoning command dedicated to breaking updates.
@@
    A large scale experiment with five different LLMs and eight prompts, over 103 breaking dependency updates. Our results show that o3-mini achieves the best outcomes when using a prompt that highlights the erroneous line causing the compilation error(s), that provides information about the API differences between the old and updated dependency versions, and that includes Chains-of-thought instructions.
@@
2 Background
Breaking dependency updates are common in software projects (Brito et al., 2018b ), and often discourage developers from upgrading their dependencies (Dietrich et al., 2019; Ochoa et al., 2022; Venturini et al., 2023). However, keeping dependencies up to date is crucial for security, as outdated libraries may contain vulnerabilities  (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020; Larios Vargas et al., 2020; Salza et al., 2018). A breaking dependency update occurs when a library update introduces incompatibilities that cause client code to fail. This can occur either when developers manually update dependencies or when they use automated tools such as Dependabot or Renovate, which submit pull requests for updates. Such breakages are often changes in method signatures, deprecations, or class removals can cascade into compilation errors across the client project. While breaking updates can also trigger test failures, dependency resolution issues or lock conflicts (Reyes et al., 2024b ), in this paper we focus specifically on compilation errors, which are the most prevalent form (Reyes et al., 2024a ).
Prior work has extensively studied the prevalence and causes of breaking changes and methods to identify them in library code (Brito et al., 2018b ; Xavier et al., 2017; Brito et al., 2020; Brito et al., 2018a ; Mujahid et al., 2020; Reyes et al., 2024a ). Migration support has also been explored through rule-based transformations and recommendation systems (Xing and Stroulia, 2007; Dagenais and Robillard, 2009; Brito et al., 2018a ). However, such approaches typically rely on predefined rules, mined examples, or static analysis heuristics, which limit their ability to handle the diversity of real-world breaking updates and the project-specific context required to repair client code. In contrast, we investigate whether large language models (LLMs) can directly generate client-side repairs when provided with appropriate context.
To formalize our scope, we adopt standard terminology.
@@ 2 Background
A breaking dependency update occurs when updating a dependency version introduces incompatibilities that cause the build to fail.
For systematic evaluation in this study, we examine breaking dependency updates as represented in the BUMP dataset Reyes et al., 2024b : a pair of commits for a project, consisting of a pre-breaking commit with a passing build and a breaking commit that updates the version of a single dependency, causing the build to fail.
@@
A breaking commit is the project commit where the only change is an update to a dependency version that leads to the build failing.
3 Byam: An Approach to Repairing Breaking Updates
Existing solutions for breaking dependency updates have limited scalability and adaptability. They require predefined rules or abundant historical data, which makes them weak in the face of diverse or novel API changes. Large language models (LLMs) offer a promising alternative due to their ability to generalize across different contexts and reason about unknown patterns. To be effective in automating the repair of breaking dependency updates, LLMs must be guided with rich prompts.
We propose a novel approach, Byam, based on generating a code patch with LLMs to fix breaking dependency updates. Byam examines error messages, locates the erroneous line, and consults API differences before querying an LLM and applying a fix. Byam enables LLMs to operate as context-aware repair assistants, making them more efficient and accurate at addressing breaking changes than prior approaches.
@@ 3.2 Step 1: Extract Failed Build Information
This step isolates the parts of the code that are causing a compilation failure after the dependency update and identifies the API differences involved in the failure. Byam uses this information to provide the LLM with various contextual data. To collect this information, Byam analyzes the build output log after a breaking update. The pipeline automatically extracts a list of files causing the compilation error(s) and the list of error messages for each file. Additionally, Byam analyzes the changes between both versions of the dependency.
Figure 1(c) shows an excerpt of a build output. Analyzing this output, Byam determines that the file /billy/billy-gin/src/main/java/com/premiumminds/billy/gin/services/impl/pdf/FOPPDFTransformer.java is causing a compilation error. Byam also extracts the error message, “no suitable method found for newInstance(no arguments), as well as the error location, at line 115, column 43. From the error location, Byam identifies the specific constructs responsible for the compilation failure. In the previous example, the error is caused by a method newInstanc(). Next, Byam extracts the API differences between the dependency versions to determine whether any changes in the new version are related to the construct that triggers the error. We term this output the API Difference (APIDiff). For this example, the APIDiff shows that the signature of the method newInstance()has been changed in the new version 2.0 of the dependency. Figure 3 shows an excerpt of the APIDiff corresponding to the example in Figure 1, highlighting the change associated with the construct identified as the root cause of the breaking dependency update. This APIDiff information is later incorporated into the LLM prompts, as discussed in Subsection 3.3.
@@
+++ NEW METHOD: PUBLIC(+) STATIC(+) org.apache.fop.apps.FopFactory newInstance(org.apache.fop.apps.FopFactoryConfig)
+++ NEW METHOD: PUBLIC(+) STATIC(+) org.apache.fop.apps.FopFactory newInstance(java.io.File)
+++ NEW EXCEPTION: org.xml.sax.SAXException
+++ NEW EXCEPTION: java.io.IOException
+++ NEW METHOD: PUBLIC(+) STATIC(+) org.apache.fop.apps.FopFactory newInstance(java.net.URI)
+++ NEW METHOD: PUBLIC(+) STATIC(+) org.apache.fop.apps.FopFactory newInstance(java.net.URI, java.io.InputStream)
+++ NEW EXCEPTION: org.xml.sax.SAXException
+++ NEW EXCEPTION: java.io.IOException
@@ 3.3 Step 2: Use LLM to Generate Repairs
Erroneous Line Inclusion Adds the specific line of code causing the compilation error.
@@
Based on the baseline, we evaluate specific combinations of prompt design choices rather than generating all possible variations, focusing on erroneous lines, API version differences (APIDiff), and Chain of Thought (CoT) reasoning. Instead of an exhaustive exploration, we systematically test selected configurations that provide meaningful information on their impact on LLM performance. Our approach considers the incremental addition of each factor to a baseline prompt, ensuring a balanced assessment of each factor’s contributions. This approach examines two key dimensions: the prompting strategy (zero-shot vs. CoT) and the type of contextual information included in the prompt (error information, erroneous line, APIDiff). This structured evaluation allows us to isolate the effects of each component while maintaining a manageable number of experimental setups.
Providing Erroneous Line Information. Developers, when debugging, typically start by identifying the exact line of code causing the issue, as this helps narrow down the scope of the problem and focus efforts on the relevant section of the code. Including the erroneous line in the prompt replicates this natural debugging process, potentially improving code understanding for the LLM (the compiler error message does not include the actual line by default). We include this information by extending the base prompt with the actual line of the client code responsible for the compilation error, as shown in Figure 5, after the error message.
@@
Figure 5: Adding Erroneous Line to the Prompt. This helps the LLM to identify and focus on the problem to be fixed.
@@
To extract these differences, we rely on japicmp (Maven, 2022), a widely used tool to detect API changes in Java libraries. japicmp performs a comprehensive structural comparison, allowing it to detect modifications across public, internal, and even deprecated parts of an API. This tool identifies modifications such as changes in method signatures, parameter types, deprecated methods, or new functionality. For example, when we use japicmp to analyze the changes between versions 6.18.1 and 6.19.1 of the jasperreports dependency, it detects that the setLinewith method was removed from the new version of the dependency. We augment the prompt with that information, as in the example shown in Figure 6. Figure 7 shows the concrete example of APIDiff for jasperreports.
@@
$P_{4}$ Erroneous Line + APIDiff
@@
$P_{8}$ CoT + Erroneous Line + APIDiff
@@ 4.1 Research Questions
In this section, we present our three research questions designed to evaluate complementary aspects of Byam’s effectiveness to repair compilation errors caused by breaking dependency updates.
@@ 1. RQ1
    This question aims to measure Byam’s capability to fully repair projects with breaking dependency updates by fixing all files and errors to reach a build success. This metric indicates how effective Byam is as a fully automated solution without developer intervention.
@@ 2. RQ2
    This question evaluates the extent to which Byam can repair projects partially, even when the build is not fully fixed. While RQ1 focuses on complete build repair, there are cases where Byam cannot resolve all errors but still fixes a substantial portion of them. Such partial repair is valuable in practice, as it reduces the number of remaining errors and allows developers to concentrate their effort on a smaller scope. In this RQ, we therefore measure partial repair at both the file level and the individual error level.
@@ 3. RQ3
    In this question, we address the balance between fixing existing errors caused by the breaking dependency update and introducing new ones when applying patches generated by Byam. Minimizing the introduction of new errors with generated fixes is essential for the practical adoption of any automated repair tool, as newly introduced errors will increase debugging effort and affect the developers’ confidence in the automated system.